OpenStudy (anonymous):

Are the following functions periodic? f(t) = Abs[t]^3 over -1/2 < t < 1/2 and f(t) = f(t+1).

OpenStudy (anonymous):

I don't see how the first function is periodic, given that for the same value for f(t), a different distance between points on the graph arises as you move up or down it.

OpenStudy (anonymous):

For the second graph, I'm not quite sure how to interpret it.

OpenStudy (anonymous):

This is all pre-fourier analysis and I know that each function needs to be continuous and their derivatives sectionally continuous. However, that still doesn't tell me that the functions are periodic which I'm having trouble discerning.

OpenStudy (sidsiddhartha):

yes those are dirichlets conditions but u dont need them to check periodicity

OpenStudy (sidsiddhartha):

for the first function u can see it looks like 

OpenStudy (anonymous):

Yeah

OpenStudy (sidsiddhartha):

so its not really periodic

OpenStudy (anonymous):

I agree. The second function is though? It looks a lot like the definition of a function that repeats itself every period, i.e. f(t) = f(t + T)

OpenStudy (sidsiddhartha):

yes its the basic definition A function f is said to be periodic with period P, if \[f(x+P)=f(x)\]

OpenStudy (sidsiddhartha):

so the second function is periodic with period 1

OpenStudy (anonymous):

Excellent. Thanks for clarifying it. Much appreciated

OpenStudy (sidsiddhartha):

no problem :)